A company wants to invest $200,000 in a piece of equipment with a 10 year service life. Additional operating costs of $5,000 per year are expected. MARR is 12%. The equipment will have a salvage value of $10,000 at the end of service.

(a) What is the annual equivalent cost of this investment? (b) If the equipment would produce 10,000 widgets per year, and the company can purchase widgets for $4.50 each, should the company make or buy the widgets?
Answer #1

a) Annual equivalent cost = 200,000(A/P, 12%, 10) + 5,000 - 10,000(A/F, 12%, 10)

                                         = 200,000(0.1770) + 5,000 - 10,000(0.0570)

                                         = 35,400 + 5,000 - 570

                                         = $39,380

b) The purchasing cost = $4.50 * 10,000 = $45,000. If the company will make widgets, it will cost $39,380. So, the company should make the widgets.